OpenStudy (anonymous):

evaluate the limit lim x->0 (5+x)^-1-5^-1/(x). The answer is -1/25 but I don't know how.

myininaya (myininaya):

\[\lim_{x \rightarrow 0}((5+x)^{-1}-5^{\frac{-1}{x}})\] correct?

OpenStudy (jdoe0001):

" The answer is 1/25" <--- seems that's found already =)

OpenStudy (anonymous):

the answer is 1/25 but I don't know how

myininaya (myininaya):

I don't think I'm seeing the question correctly, could be wrong? But I don't the above I think I wrote has a limit of 1/25.

myininaya (myininaya):

\[\lim_{x \rightarrow 0}\frac{(5+x)^{-1}-5^{-1}}{x}\]?

OpenStudy (anonymous):

yes

myininaya (myininaya):

do you know derivatives because you can skip a lot of steps if you do

OpenStudy (anonymous):

I do know them

myininaya (myininaya):

\[\lim_{x \rightarrow a}\frac{f(x)-f(a)}{x-a}=f'(x)|_{x=a}\] are you familiar with this definition?

OpenStudy (anonymous):

sort of, but I'm more familiar with f(x+h)-f(x)/h

myininaya (myininaya):

\[\lim_{x \rightarrow 0}\frac{(5+x)^{-1}-(5+0)^{-1}}{x-0}=\lim_{x \rightarrow a}\frac{f(x)-f(a)}{x-a}\] If you are good at comparing you will be able to see f(x)=(5+x)^(-1) and a=0

myininaya (myininaya):

\[\lim_{x \rightarrow 0}\frac{(5+x)^{-1}-(5+0)^{-1}}{x-0}=[(5+x)^{-1}]'|_{x=0}\]

myininaya (myininaya):

so you will have to know chain rule to do it this way.

myininaya (myininaya):

If you don't like this way we can do it the good old fashion way.

OpenStudy (anonymous):

yes please, if its not any trouble

OpenStudy (anonymous):

and the answer is -1/25, sorry

myininaya (myininaya):

\[\lim_{x \rightarrow 0}\frac{\frac{1}{5+x}-\frac{1}{5}}{x}=\lim_{x \rightarrow 0}\frac{1}{x}(\frac{1}{x+5}-\frac{1}{5})\] We cannot just plug in 0 because the bottom will be 0. We want to get rid of that factor of x on the bottom. So sometimes a trick when seeing separate fractions is to combine them.

myininaya (myininaya):

An easy way to combine fractions is: